# thanks: Corresponding authorthanks: left academia

### v1

@@ V Momentum relaxation and equilibration time
to the obtained time evolution curves of Fig. 8. The resulting equilibration times—together with their respective relaxation times—are shown in Table 1.
@@
Table 1: Comparison of the relaxation and equilibration times for each of the curves in Fig. 8 with different scaling factor of the drag force coefficient ($k\gamma$) used in the simulation at $T=160$ MeV.
@@ VI Equilibrium Yields and Comparison to Statistical Hadronization Model
where the gluons represents the random kicks from the system which the charm and anticharm quarks experience. In this reaction, the chemical potentials of the reactives and the product are connected through the relation $\mu_{c\bar{c}}=2\mu_{c}$, leading to $\lambda_{c\bar{c}}=\lambda_{c}^{2}$ for the fugacity.
@@ VII Conclusions and Outlook
After having passed all the tests within box simulations and ensuring, that the correct equilibrium limit is reached, and the bound state formation, dissociation and regeneration occurs in the expected manner we can now move forward and adapt our model to a situation of a realistic heavy-ion collision. In the future we will embed our model in an expanding medium, modeled by a blast-wave like elliptic fireball. Within this dynamical description the heavy quarks will initially be placed according to the Glauber model and with an initial momentum distribution given by PYTHIA. Using this setup, we can now furthermore study the elliptic flow of charm quarks and charmonia as well as the nuclear modification factor at RHIC and LHC energies. Since up to this point only regenerated charmonium states have been included in the description, it should be considered to additionally include primordial $J/\psi$ to be able to better compare to experimental data. This is especially important at RHIC energies, where regeneration processes are believed to be negligible and the dominant effect is the dissociation and suppression of initial bound states. This is work in progress. Furthermore, the description allows to be expanded to the bottom-sector and the bottomonium formation, as well as $B_{c}$ states.

### v3

@@ V Momentum relaxation and chemical equilibration time
to the obtained time evolution curves of Fig. 9. The resulting chemical equilibration times—together with their respective relaxation times—are shown in Table 1.
@@
Table 1: Comparison of the relaxation and chemical equilibration times for each of the curves in Fig. 9 with different scaling factor of the drag force coefficient ($k\gamma$) used in the simulation at $T=160$ MeV.
@@ VI Equilibrium Yields and Comparison to Statistical Hadronization Model
where the gluons represents the random kicks from the system which the charm and anticharm quarks experience. In this reaction, the chemical potentials of the reactive and the product are connected through the relation $\mu_{c\bar{c}}=2\mu_{c}$, leading to $\lambda_{c\bar{c}}=\lambda_{c}^{2}$ for the fugacity.
@@ VII Conclusions and Outlook
After having passed all the tests within box simulations and ensuring, that the correct chemical equilibrium limit is reached, and the bound state formation, dissociation and regeneration occurs in the expected manner we can now move forward and adapt our model to a situation of a realistic heavy-ion collision. In the future we will embed our model in an expanding medium, modeled by a blast-wave like elliptic fireball. Within this dynamical description the heavy quarks will initially be placed according to the Glauber model and with an initial momentum distribution given by PYTHIA. Using this setup, we can now furthermore study the elliptic flow of charm quarks and charmonia as well as the nuclear modification factor at RHIC and LHC energies. Since up to this point only regenerated charmonium states have been included in the description, it should be considered to additionally include primordial $J/\psi$ to be able to better compare to experimental data. This is especially important at RHIC energies, where regeneration processes are believed to be negligible and the dominant effect is the dissociation and suppression of initial bound states. This is work in progress. Furthermore, this description allows to be expanded to the bottom sector and the bottomonium formation, as well as $B_{c}$ states. Finally, since the complete description of a heavy-ion collision should also account for the hadronic medium evolution Das et al. (2024), the whole framework could also be applied to the confined phase to describe $D$ and $\bar{D}$ mesons in a bath of light constituents, with the possibility of the $X(3872)$ formation or other molecular states at finite temperature Montaña et al. (2023).
@@ Acknowledgments
We thank Taesoo Song for valuable discussions on the topic and for providing to us the DQPM data presented in the lower panel of Fig. 2.